# One-Shot Real-to-Sim via End-to-End Differentiable Simulation and Rendering

### v3

@@ Abstract
Identifying predictive world models for robots in novel environments from sparse online observations is essential for robot task planning and execution in novel environments. However, existing methods that leverage differentiable programming to identify world models are incapable of jointly optimizing the geometry, appearance, and physical properties of the scene. In this work, we introduce a novel rigid object representation that allows the joint identification of these properties. Our method employs a novel differentiable point-based geometry representation coupled with a grid-based appearance field, which allows differentiable object collision detection and rendering. Combined with a differentiable physical simulator, we achieve end-to-end optimization of world models, given the sparse visual and tactile observations of a physical motion sequence. Through a series of world model identification tasks in simulated and real environments, we show that our method can learn both simulation- and rendering-ready world models from only one robot action sequence.
@@ I Introduction
Many existing works take advantage of differentiable simulators as strong physics priors that allow efficient identification of physical parameters such as inertia and coefficient of friction [5, 6, 7]. Differentiable simulators allow the gradient-based optimization of mass-inertial properties and frictional coefficients to match a physical motion sequence to sparse robot observations. However, these works assume known geometries and appearances of the objects in the scene and do not allow the algorithm to adapt the GAP simultaneously.
@@
To address these challenges, this work presents an object representation that is compatible with general-purpose rigid body simulators and allows the joint optimization of GAP. As shown in Fig. 1, based on this representation, our work enables the identification of a rigid world model in the form of a full-fledged rigid body simulator from the observations of one robot push. Our proposed representation is the combination of a recently proposed point-based shape representation Shape-as-Points (SaP) [11] and a grid-based appearance field. SaP parameterizes an object’s geometry and topology using a set of surface points along with normal directions. It then uses differentiable Poisson reconstruction to recover a smooth indicator field of object occupancy, which can be converted to a mesh using a differentiable marching cubes algorithm [12]. The texture of the vertices of the mesh is then obtained by interpolating the appearance grid. Employing the mesh in a differentiable rigid body simulator [13] that provides gradients for the physical parameters and contact points of the objects, our method constructs a fully differentiable pipeline for jointly optimizing the geometry, appearance, and physical parameters. Our contributions are:
  • A jointly differentiable representation of the shape, appearance, and physical properties of objects.
@@
We evaluate our method on a series of identification problems in both simulated and real-world environments. The results show that our method can infer accurate world models from a single episode of robot interactions with the environment. Finally, we will release the code and data for reproducing the experiments in this paper once the manuscript is accepted.
@@ II-C World Models
Traditional system identification methods [24] identify only the dynamics parameters from full state information. However, to support diverse downstream robot tasks in the real world, world models need to be built from raw observations and support both accurate dynamics prediction and photorealistic novel view synthesis. While existing works have identified world models from raw image observations using differentiable simulators [17, 13, 15], none supports simultaneous optimization of geometry, appearance, and physical parameters. Another line of work closely related to ours is image-based generative world modeling. These works aim to predict the next RGB frame based on the current frame and action. These models are learned by training on diverse datasets with generative modeling techniques such as variational autoencoders [1, 2] and diffusion [3, 4]. The key differences between our method and these works are that our simulation, grounded in physics, is always physically consistent and is a general-purpose rigid simulator that can provide physical information such as contact forces. Purely data-driven world models generalize poorly to novel scenarios and their lack of physical information severely limits their application to downstream robot tasks. Finally, a recent work [25] proposed a method to use Gaussian Splatting along with a particle-based simulator to track and reconstruct a moving scene. Instead of identifying the physical parameters of the scene, the method optimizes virtual forces attached to each particle such that they match the observed object trajectory. Therefore, although the method can be used as a world model for prediction, the accuracy is severely limited. We include this method as a baseline in our experiments in Sec. V and demonstrate the limitation of this method.
@@ III Problem Definition
In this section, we describe our formulation of world model identification. We assume the environment consists of a rigid object and rigid terrain, whose physical properties and appearances are parameterized by $\theta$. A robot, equipped with joint encoders and end-effector force sensors, interacts with the object at $T$ time instances: $t_{1},\cdots,t_{T}$, with a fixed time step $\delta t$. At each time step, the robot observes its end-effector pose $e^{t}\in\text{SE}(3)$ and contact force $f^{t}\in\mathbb{R}^{3}$. Further, the robot is equipped with an RGB-D camera with known intrinsics that observes the object through image $o^{t}\in\mathbb{R}^{H\times W\times 4}$ at camera pose $c^{t}\in\text{SE}(3)$. We further assume an image segmentation mask $m^{t}\in\mathbb{R}^{H\times W\times 4}$ is provided for the object, robot, and terrain. Therefore, the robot observations are a sequence $\mathcal{O}=\{\langle t,e^{t},f^{t},o^{t},c^{t},m^{t}\rangle\}$, and our goal is to estimate $\theta$ from the set of sparse observations $\mathcal{O}$. We formulate this problem as a physics-constrained optimization by introducing a full-fledged physics simulator function $q^{i+1}=g(q^{i},u^{i},\theta)$ that can differentiate through objects’ appearance, geometry, and physical parameters. Here $q^{i}\in\text{SE}(3)\times\text{SE}(3)$ is the object and robot end-effector poses at timestep $t$ and $u^{i}$ is the applied robot force at the end-effector, which is equal in magnitude to the sensed contact force but opposite in direction.
@@
$\mathbf{s.t.}$ $q^{i+1}=g(q^{i},u^{i},\theta)\quad\forall i=1,\cdots,T.$
over a physical motion sequence of $T$ timesteps, with the objective function $\mathcal{L}$ encourages the simulated observation $\hat{o}^{t}$ to match the ground-truth observation $o^{t}$.
@@ IV-A Differentiable Object Representation
where the first two parameters are mass-inertial properties and frictional coefficients, defining the physical properties, and the last two parameters are the oriented point cloud for SaP and color field, defining the objects’ appearance.
@@ IV-B Differentiable Simulator
The governing equation of motion for rigid bodies and the time discretization method is well-studied and we refer the readers to Anitescu et al. [31] for details. The equation is summarized as follows:
@@
Figure 3: The experiment setups for the simulated (left) and physical (right) experiments. 9 objects are used for simulation with the PyBullet simulator, including 8 YCB objects and a green box. For the real-world experiments, three YCB objects (Drill, Mustard, and Sugar) are used. A UR5e arm equipped with a pusher and an overhead Realsense D435 RGB-D camera are used. Note that only the circled object in the real-world setup is the object of interest and everything else is treated as the static terrain.
@@ IV-C World Model Identification
Even with our jointly differentiable physical and appearance models, solving Eqn. 1 can still be rather challenging. This is mainly because our initial guess can be very poor, especially in the occluded region. As a result, the naïve gradient descent method can take many iterations and is prone to converging to poor local minima. To mitigate this, we use two stages of optimization, and we further leverage 3D foundation models trained on web-scale data to generate good initial guesses of the rigid object in the scene from partial visual observations.
@@ IV-C1 Two-stage Optimization
with $(c_{1},\dots,c_{7})$ denoting weight terms. Here, $\hat{o}^{t_{1}}$ is the rendered RGB-D image at $t_{1}$. $\mathcal{L}_{\text{rgb}}$ is a loss on the RGB images, defined as a weighted sum of $l_{1}$ distance and D-SSIM terms: $\mathcal{L}_{\text{rgb}}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\text{SSIM}}$, where we set $\lambda=0.2$. $\mathcal{L}_{\text{depth}}$ is the $l_{1}$ distance on the depth images. When calculating the three loss terms $\mathcal{L}_{1,\text{SSIM},\text{depth}}$, the robot is masked out according to the segmentation mask $m^{t_{1}}$. $\mathcal{L}_{\text{pcd}}$ is a unilateral Chamfer distance between the point cloud generated from the observed RGB-D image pixels belonging to the object and the SaP point cloud $\mathcal{P}$, which is defined as the average of the minimum distance between each point on the observed point cloud and any point among the SaP points. The terms $\mathcal{L}_{\text{pen}}$ and $\mathcal{L}_{\text{balance}}$ encourage the object to achieve static force equilibrium while having no penetrations. $\mathcal{L}_{\text{pen}}$ penalizes the object mesh penetrations in the terrain, and $\mathcal{L}_{\text{balance}}=\sum_{1}^{k}\|p_{o}^{i+t_{1}}-p_{o}^{t_{1}}\|^{2}$ is the sum of the positional changes of the object from the initial position $p_{o}^{t_{1}}$ over $k$ steps by simulating forward with no robot actions. In the case where the initial guess of the object geometry does not come in contact with the terrain at $t_{1}$, $\mathcal{L}_{\text{balance}}$ allows the computation of the gradient information to expand the geometry towards the terrain once the object falls due to gravity and contacts the terrain during the $k$ steps. We use $k=3$ for all the experiments in this paper. These two terms provide a strong hint for the occluded part of the object. For example, when an object is lying on the table, our RGB-D observation will not cover the bottom of the object. However, our model will guide SaP to fill the bottom-side geometries by encouraging the object to settle on the table. Finally, the last two terms regularize the object shape, where ${L}_{\text{reg}}$ is the $L_{2}$ norm between the SaP points and the initial SaP points and $L_{smooth}$ is the Laplacian smoothing objective on the object mesh.
@@
In addition, the use of both $\mathcal{L}_{\text{depth}}$ and $\mathcal{L}_{\text{pcd}}$ is necesary. When only $\mathcal{L}_{\text{depth}}$ is used, if the estimated mesh is smaller than the ground-truth object geometry, the predicted depth pixels that are supposed to reach the ground-truth mesh do not hit the estimated mesh, and there is no gradient information for expanding the geometry. Similarly, the observed point cloud used in $\mathcal{L}_{\text{pcd}}$ is based on a partial point cloud of the object, and SaP has no information about the boundaries of the object. $\mathcal{L}_{\text{depth}}$ informs our model of the background area and prevents the object geometries from occupying the supposed background.
@@
After the first stage, we have tuned our model to match the first observation. When we move on to the second stage, we incorporate all timesteps by applying the robot controls. For this stage, we assume the current estimated geometry is close to the ground truth and use the simpler loss: $\mathcal{L}=c_{8}\mathcal{L}_{\text{pcd}}+c_{9}\mathcal{L}_{\text{robot}}$, where the first term has the same definition as in the first stage and $\mathcal{L}_{\text{robot}}$ is the squared distance between the ground-truth and predicted robot joint angles. Instead of using the observations of every frame for calculating the loss that would be computationally very expensive, we apply the robot control forces from the initial state, integrate forward in time, and render two intermediate and the last frames for calculating the losses. During our optimization, the chain of gradient is back-propagated through the following recursive rule:
@@ IV-C2 Geometry Prior
Our method relies on a reasonable initial guess. Imagine the case with an object settling on the edge of a table and the camera does not observe the contact between the two. The initial guess of the occluded part of the object could be very short and cause the object to directly fall down without touching the table. This cannot be recovered by our optimization since the object never hits the table and there are no gradients for correcting the geometries. To obtain a reasonable initial guess of the geometries and appearance of the rigid object of interest from partial visual observations, we take advantage of large reconstruction models [34] that predict object 3D models from a single RGB image, trained on web-scale data. In particular, we use TripoSR [35] in our experiments with the segmented RGB image of the object as the input image. Since the generated mesh is scale- and transform-agnostic, we subsequently apply RANSAC and scale-aware iterative closest point (ICP) algorithms with Open3D [36] to register the mesh to the partial object point cloud, computed from the RGB-D image at the first time instance.
@@ V Experiments and Results
Figure 7: Results of three example testing trajectory of the physical experiments. The predicted poses are obtained by applying the control forces from the initial pose and integrating forward in time. The predicted object poses with the optimized $\theta$ highlighted with a yellow silhouette are overlaid with the ground-truth object and robot, and the background rendered from the optimized simulator. [Best viewed in color.]
@@ V-A Simulation Experiments
Shown in Fig. 3, we conduct all the simulated experiments using data collected with the PyBullet simulator [37]. We use a simple green box object (Box) and 8 objects (Gelatin, RubiksCube, Spam, TunaCan, Mustard, Bleach, Drill, and Sugar) from the YCB object dataset [38]. The objects are placed on a flat surface with checker patterns and pushed by a floating sphere robot, while a static overhead camera takes pictures. The world model parameters are first optimized with data collected on one training push, and evaluated on 23 novel pushes that are drastically different. Fig. 4 shows the starting positions and the pushing directions of the pushes for the Drill object. Similar pushes are used for all the other objects where the starting locations are adjusted based on the size of the objects. The trajectories push the objects up to 12 cm and 80. All the trajectories have $T=30$ time steps with $\delta t=0.01$ s. We use the following weights for optimization: $[c_{1},c_{2},c_{3},c_{4},c_{5},c_{6},c_{7},c_{8},c_{9}]=[10,5000,500,1,1000,100,4000,500,100]$. These terms are not carefully tuned and are set such that each term has a similar order of magnitude at the start of optimization for our experiments. The physical parameters we optimize for include the mass, surface coefficient of friction, the center of mass, and the rotational inertial. The center of mass is initialized at the geometry center of the initial shape guess. The rotational inertia is initialized by treating the object as a box, whose dimensions are the bounding box of the initial geometry guess. We assume that the rotational inertia only has diagonal terms. For all experiments including simulations and real-world experiments, the surface coefficient of friction is initialized at 0.2 and the mass is initialized at 0.2 kg.
@@
We first evaluate whether our method can identify the shape and physical parameters accurately such that it generalizes well to new physical interactions. We compare our method against a recently proposed method that represents the world jointly with Gaussian Splats and physical particles and allows it to perform both novel-view rendering and physics-based trajectory predictions. We refer to this method as PhysGS. This method performs simulation with a particle-based simulator [39]. In the original paper, the physical parameters of the particles are arbitrarily set, and virtual forces are optimized to match the observations and predictions on the training trajectory. To allow more accurate prediction for new trajectories, we optimize the total mass and coefficient of friction of the object particles with grid search using the partial Chamfer distance between the observed object point cloud and the predicted object particles on the last time frame in the trajectory.
@@
We report the quantitative results for dynamics parameter estimation and novel trajectory predictions for all objects in the simulation experiments in Table I, which are the average across all objects. We also shown some qualitative examples that are representative of the average errors in Fig. 5. The dynamics parameter estimation error from the training trajectory, and the average pose error, unilateral chamfer distance, and the translational velocity error at the end of the testing trajectories are reported. The final pose and velocity of the objects are obtained by applying the control forces and integrating forward in time. Our method identifies the dynamics parameters accurately and shows low trajectory prediction errors. We would like to point out in particular that the average rotation error is heavily skewed by the TunaCan object since our method currently uses the object surface point cloud to track the object during stage 2 optimization, and cannot properly differentiate the rotation of a cylindrical object. As a result, the average rotation error is 28.0 for TunaCan, and we aim to address this issue in future work. On the other hand, PhysGS generalizes very poorly to the testing trajectories. While this is partially because of the lack of proper dynamics parameters and physics-based shape estimation, we also find that the particle-based simulator is extremely sensitive to simulator parameters and have poor physical fidelity, especially for rigid objects. We also show an example of the initial and optimized geometries of the Box object in Fig. 8 (left), which demonstrate the ability of our method to adjust occluded geometry based on the physics.
Next, we also evaluate the quality of the novel-view synthesis of our method. For each of the testing objects, we evaluate the synthesized RGB images from 10 novel viewpoints around the scene, and our method achieves 0.00225 of mean squared error (MSE), 0.965 of structural similarity index measure (SSIM) and 26.5 of peak signal-to-noise ratio (PSNR). We also show some examples of novel view synthesis of the Drill object in Fig. 6, which matches the ground truth very well.

### v4

@@ Abstract
Identifying predictive world models for robots from sparse online observations is essential for robot task planning and execution in novel environments. However, existing methods that leverage differentiable programming to identify world models are incapable of jointly optimizing the geometry, appearance, and physical properties of the scene. In this work, we introduce a novel rigid object representation that allows the joint identification of these properties. Our method employs a novel differentiable point-based geometry representation coupled with a grid-based appearance field, which allows differentiable object collision detection and rendering. Combined with a differentiable physical simulator, we achieve end-to-end optimization of world models or rigid objects, given the sparse visual and tactile observations of a physical motion sequence. Through a series of world model identification tasks in simulated and real environments, we show that our method can learn both simulation- and rendering-ready rigid world models from only one robot action sequence. The code and additional videos are available at our project website: https://tianyi20.github.io/rigid-world-model.github.io/.
@@ I Introduction
Many existing works employ differentiable simulators as strong physics priors that allow efficient identification of physical parameters such as inertia and coefficient of friction [5, 6, 7]. Differentiable simulators allow the gradient-based optimization of mass-inertial properties and frictional coefficients to match a physical motion sequence to sparse robot observations. However, these works assume known geometries and appearances of the objects in the scene and do not allow the algorithm to adapt the GAP simultaneously.
@@
To address these challenges, this work presents a rigid object representation that is compatible with general-purpose rigid body simulators and allows the joint optimization of GAP. As shown in Fig. 1, based on this representation, our work enables the identification of a rigid world model in the form of a full-fledged rigid body simulator from the observations of one robot push. Our proposed representation is the combination of a recently proposed point-based shape representation Shape-as-Points (SaP) [11] and a grid-based appearance field. SaP parameterizes an object’s geometry and topology using a set of surface points along with normal directions. It then uses differentiable Poisson reconstruction to recover a smooth indicator field of object occupancy, which can be converted to a mesh using a differentiable marching cubes algorithm [12]. The texture of the vertices of the mesh is then obtained by interpolating the appearance grid. Employing the mesh in a differentiable rigid body simulator [13] that provides gradients for the physical parameters and contact points of the objects, our method constructs a fully differentiable pipeline for jointly optimizing the GAP. Our contributions are:
  • A jointly differentiable representation of the shape, appearance, and physical properties of rigid objects.
@@
We evaluate our method on identification problems in both simulated and real-world environments. The results show that our method can infer accurate world models from a single episode of robot interactions with the environment.
@@ II-C World Models
Traditional system identification methods [24] identify only the dynamics parameters from full state information. However, to support diverse downstream robot tasks in the real world, world models need to be built from raw observations and support both accurate dynamics prediction and photorealistic novel view synthesis. While existing works have identified world models from raw image observations using differentiable simulators [17, 13, 15], none supports simultaneous optimization of GAP. Another line of work closely related to ours is image-based generative world modeling. These works aim to predict the next RGB frame based on the current frame and action. These models are learned by training on diverse datasets with generative modeling techniques such as variational autoencoders [1, 2] and diffusion [3, 4]. The key differences between our method and these works are that our simulation, grounded in physics, is always physically consistent and is a general-purpose rigid simulator that can provide physical information such as contact forces. Purely data-driven world models generalize poorly to novel scenarios and their lack of physical information severely limits their application to downstream robot tasks. Finally, a recent work [25] proposed a method to use Gaussian Splatting along with a particle-based simulator to track and reconstruct a moving scene. Instead of identifying the physical parameters of the scene, the method optimizes virtual forces attached to each particle such that they match the observed object trajectory. Therefore, although the method can be used as a world model for prediction, the accuracy is severely limited. We include this method as a baseline in our experiments in Sec. V and demonstrate the limitation of this method.
@@ III Problem Definition
In this section, we describe our formulation of world model identification. We assume the environment consists of a rigid object and rigid terrain, whose physical properties and appearances are parameterized by $\theta$. A robot, equipped with joint encoders and end-effector force sensors, interacts with the object at $T$ time instances: $t_{1},\cdots,t_{T}$, with a fixed time step $\delta t$. At each time step, the robot observes its end-effector pose $e^{t}\in\text{SE}(3)$ and contact force $f^{t}\in\mathbb{R}^{3}$. Further, the robot is equipped with an RGB-D camera with known intrinsics that observes the object through image $o^{t}\in\mathbb{R}^{H\times W\times 4}$ at camera pose $c^{t}\in\text{SE}(3)$. We further assume an image segmentation mask $m^{t}\in\mathbb{R}^{H\times W\times 4}$ is provided for the object, robot, and terrain. Therefore, the robot observations are a sequence $\mathcal{O}=\{\langle t,e^{t},f^{t},o^{t},c^{t},m^{t}\rangle\}$, and our goal is to estimate $\theta$ from the set of sparse observations $\mathcal{O}$. We formulate this problem as a physics-constrained optimization by introducing a full-fledged physics simulator function $q^{i+1},\dot{q}^{i+1}=g(q^{i},\dot{q}^{i},u^{i},\theta)$ that can differentiate through objects’ appearance, geometry, and physical parameters. Here, $q^{t}$ and $\dot{q}^{i+1}$ are the object and robot end-effector poses and velocities at timestep $i$ and $u^{i}$ is the applied robot force at the end-effector, which is equal in magnitude to the sensed contact force but opposite in direction.
@@
$\mathbf{s.t.}$ $q^{i+1},\dot{q}^{i+1}=g(q^{i},\dot{q}^{i},u^{i},\theta)\quad\forall t=1,\cdots,T-1.$
The optimization is solved over a physical motion sequence of $T$ timesteps, with the objective function $\mathcal{L}$ encourages the simulated observation $\hat{o}^{t}\big{(}q^{t}(\theta),\theta\big{)}$ to match the ground-truth observation $o^{t}$.
@@ IV-A Differentiable Object Representation
where the first two parameters are mass-inertial properties and frictional coefficients, and the last two parameters are the oriented point cloud for SaP and color field.
@@ IV-B Differentiable Simulator
The governing equation of motion for rigid bodies and the time discretization method are well-established, and we refer the readers to Anitescu et al. [31] for details. The equation is summarized as follows:
@@
Figure 3: The experiment setups for the simulation (left) and physical (right) experiments. 9 objects are used for simulation with the PyBullet simulator, including 8 YCB objects and a green box. For the real-world experiments, three YCB objects (Drill, Mustard, and Sugar) are used. A UR5e arm equipped with a pusher and a ATI Gamma F/T sensor and an overhead Realsense D435 RGB-D camera are used. Note that only the circled object in the real-world setup is the object of interest and everything else is treated as the static terrain.
@@ IV-C World Model Identification
Even with our jointly differentiable physical and appearance models, solving Eqn. 1 can still be rather challenging. This is mainly because our initial guess can be very poor, especially in the occluded region. As a result, the naïve gradient descent method can take many iterations and is prone to converging to poor local minima. To mitigate this, we use two stages of optimization, and we further leverage 3D foundation models trained on web-scale data to generate reasonable initial guesses of the rigid object in the scene from partial visual observations.
@@ IV-C1 Two-stage Optimization
with $(c_{1},\dots,c_{7})$ denoting weight terms. Here, $\hat{o}^{t_{1}}$ is the rendered RGB-D image at $t_{1}$. $\mathcal{L}_{\text{rgb}}$ is a loss on the RGB images, defined as a weighted sum of $l_{1}$ distance and D-SSIM terms: $\mathcal{L}_{\text{rgb}}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\text{SSIM}}$, where we set $\lambda=0.2$. $\mathcal{L}_{\text{depth}}$ is the $l_{1}$ distance on the depth images. When calculating the three loss terms $\mathcal{L}_{1,\text{SSIM},\text{depth}}$, the robot is masked out according to the segmentation mask $m^{t_{1}}$. $\mathcal{L}_{\text{pcd}}$ is a unilateral Chamfer distance between the point cloud generated from the observed RGB-D image pixels belonging to the object and the mesh vertices generated from $\mathcal{P}$, which is defined as the average of the minimum distance between each point on the observed point cloud and any mesh vertex. The use of a unilateral Chamfer distance as opposed to the regular Chamfer distance is necessary since the observed object point cloud only includes points on the object surfaces visible from the camera viewpoint. The terms $\mathcal{L}_{\text{pen}}$ and $\mathcal{L}_{\text{balance}}$ encourage the object to achieve static force equilibrium while having no penetrations. $\mathcal{L}_{\text{pen}}$ penalizes the object mesh penetrations in the terrain, and $\mathcal{L}_{\text{balance}}=\sum_{1}^{k}\|p_{o}^{i+t_{1}}-p_{o}^{t_{1}}\|_{1}$ is the sum of the positional changes of the object from the initial position $p_{o}^{t_{1}}$ over $k$ steps by simulating forward with no robot actions. In the case where the initial guess of the object geometry does not come in contact with the terrain at $t_{1}$, $\mathcal{L}_{\text{balance}}$ allows the computation of the gradient information to expand the geometry towards the terrain once the object falls due to gravity and contacts the terrain during the $k$ steps. We use $k=3$ for all the experiments in this paper. These two terms provide a strong hint for the occluded part of the object. For example, when an object is lying on the table, our RGB-D observation will not cover the bottom of the object. However, our model will guide SaP to fill the bottom-side geometries by encouraging the object to settle on the table. Finally, the last two terms regularize the object shape, where ${L}_{\text{reg}}$ is the $L_{2}$ norm between the SaP points and the initial SaP points and $L_{smooth}$ is the Laplacian smoothing objective on the object mesh.
@@
In addition, the use of both $\mathcal{L}_{\text{depth}}$ and $\mathcal{L}_{\text{pcd}}$ is necessary. When only $\mathcal{L}_{\text{depth}}$ is used, if the estimated mesh is smaller than the ground-truth object geometry, the predicted depth pixels that are supposed to reach the ground-truth mesh do not hit the estimated mesh, and there is no gradient information for expanding the geometry. Similarly, $\mathcal{L}_{\text{pcd}}$ does not inform the SaP to not expand over the observed point cloud, and $\mathcal{L}_{\text{depth}}$ prevents the object geometries from occupying the supposed background.
@@
After the first stage, we have tuned our model to match the first observation. When we move on to the second stage, we incorporate all timesteps by applying the robot controls. To make sure the geometry and appearance of the object stay close to that in the first time step, we still use the loss from the first stage on the first time frame, except for the penetration and balance losses. For the rest of the time frames, we use the loss $\mathcal{L}=c_{8}\mathcal{L}_{\text{pcd}}+c_{9}\mathcal{L}_{\text{robot}}$, where the first term has the same definition as in the first stage and $\mathcal{L}_{\text{robot}}$ is the squared distance between the ground-truth and predicted robot end-effector positions. To calculate this loss, we apply the robot control forces from the initial state, integrate forward in time, and render two intermediate and the last frames, instead of every time frame for computational efficiency. During our optimization, the chain of gradients is back-propagated through the following recursive rule:
@@ IV-C2 Geometry Prior
Our method relies on a reasonable initial guess. Imagine the case with an object settling on the edge of a table and the camera does not observe the contact between the two. The initial guess of the occluded part of the object could be very short and cause the object to directly fall down without touching the table. This cannot be recovered by our optimization since the object never hits the table and there are no gradients for correcting the geometries. To obtain a reasonable initial guess of the geometries and appearance of the rigid object of interest from partial visual observations, we take advantage of large reconstruction models [34] that predict object 3D models from a single RGB image, trained on web-scale data. In particular, we use TripoSR [35] in our experiments with the segmented RGB image of the object as the input image. Since the generated mesh is scale- and transform-agnostic, we apply RANSAC and the scale-aware iterative closest point algorithms with Open3D [36] to register the mesh to the partial object point cloud, computed from the RGB-D image at the first time instance.
@@ V Experiments and Results
Figure 7: Results of three example testing trajectory of the physical experiments. The predicted object and robot poses with the optimized $\theta$ highlighted with a yellow silhouette are overlaid with the ground-truth object, robot, and background. [Best viewed in color.]
@@ V-A Simulation Experiments
Shown in Fig. 3, we conduct all the simulated experiments using data collected with the PyBullet simulator [37]. We use a simple green box object (Box) and 8 objects (Gelatin, RubiksCube, Spam, TunaCan, Mustard, Bleach, Drill, and Sugar) from the YCB object dataset [38], which covers diverse shapes, sizes, and textures. The objects are placed on a flat surface with checker patterns and pushed by a floating sphere robot, while a static overhead camera takes pictures. As shown in Fig. 4, 24 pushing trajectories are adopted, where one is used to optimize the world model and the rest for evaluating the optimized model. To make sure that the pushes are diverse, we pick 2 starting locations on each of the four sides of the object, and push in three directions that are 20 apart at each of these 8 locations toward the object. Similar pushes are used for all the other objects where the starting locations are adjusted based on the size of the objects. The trajectories push the objects up to 12 cm and 80. All the trajectories have $T=30$ time steps with $\delta t=0.01$ s. We use the following weights for optimization: $[c_{1},c_{2},c_{3},c_{4},c_{5},c_{6},c_{7},c_{8},c_{9}]=[10,500,2000,100,100,100,4000,500,100]$. These terms are not carefully tuned and are set such that each term has a similar order of magnitude at the start of optimization for our experiments. The physical parameters we optimize for include the mass, surface coefficient of friction, the center of mass, and the rotational inertial. The center of mass is initialized at the geometry center of the initial shape guess. The rotational inertia is initialized by treating the object as a box, whose dimensions are the bounding box of the initial geometry guess. We assume that the rotational inertia only has diagonal terms. For all experiments including simulations and real-world experiments, the surface coefficient of friction is initialized at 0.2 and the mass is initialized at 0.2 kg.
@@
We first evaluate whether our method can identify the shape and physical parameters accurately, such that it generalizes well to new physical interactions. We compare our method against a recently proposed method that represents the world jointly with Gaussian splats and physical particles and allows it to perform both novel-view rendering and physics-based trajectory predictions [25]. We refer to this method as PhysGS. This method performs simulation with a particle-based simulator [39]. In the original paper, the physical parameters of the particles are arbitrarily set, and virtual forces are optimized to match the observations and predictions on the training trajectory. To allow more accurate prediction for new trajectories, we optimize the total mass and coefficient of friction of the object particles with grid search using the partial Chamfer distance between the observed object point cloud and the predicted object particles on the last time frame in the trajectory.
@@
We report the quantitative results for dynamics parameter estimation and novel trajectory predictions for all objects in the simulation experiments in Table I, which are the average across all objects. The dynamics parameter estimation error from the training trajectory, the average pose error, the unilateral Chamfer distance, and the translational velocity error at the end of the testing trajectories are reported. We also show some qualitative examples that are representative of the average errors in Fig. 5. The final pose and velocity of the objects are obtained by applying the control forces and integrating forward in time. Our method identifies the dynamics parameters accurately and shows low trajectory prediction errors. We would like to point out in particular that the average rotation error is heavily skewed by the TunaCan object since our method currently uses the object surface point cloud to track the object during stage 2 optimization, and cannot properly differentiate the rotation of a cylindrical object. As a result, the average rotation error is 28.0 for TunaCan, and we aim to address this issue in future work. On the other hand, PhysGS generalizes very poorly to the testing trajectories. While this is partially because of the lack of proper dynamics parameters and physics-based shape estimation, we also find that the particle-based simulator is extremely sensitive to simulator parameters and have poor physical fidelity, especially for rigid objects. We also show an example of the initial and optimized geometries of the Box object in Fig. 8 (left), which demonstrate the ability of our method to adjust occluded geometry based on the physics. Next, we also evaluate the quality of the novel-view synthesis of our method. For each of the testing objects, we evaluate the synthesized RGB images from 10 novel viewpoints around the scene, and our method achieves 0.00225 of mean squared error (MSE), 0.965 of structural similarity index measure (SSIM) and 26.5 of peak signal-to-noise ratio (PSNR). We also show some examples of novel view synthesis of the Drill object in Fig. 6, which matches the ground truth very well.